Find the equation for the parabola with vertex (0, 0) and directrix x = 4
Answer by lwsshak3(11628)

Find the equation for the parabola with vertex (0, 0) and directrix x = 4
***
The directrix, x=4, is a straight vertical line so the parabola opens rightwards with a horizontal axis of symmetry. Standard form of its equation: (y-k)^2=4p(x-h)^2, with (h,k) being the (x,y) coordinates of the vertex.
..
For given parabola:
p=4 (distance from vertex to directrix on axis of symmetry)
Equation: (y-0)^2=4*4(x-0)^2
Equation: y^2=16x